OpenStudy (anonymous):

A Jaguar in an auto accident in England in 1960 left the longest recorded skid marks on a public road: an incredible 290-m long. As we will see later, the friction force between the tires and the pavement varies with speed, producing a deceleration that increases as the speed decreases. Assuming an average acceleration of - 4.3 m/s2 (that is, - 0.44 g), calculate the Jag's speed when the brakes locked.

OpenStudy (anonymous):

v^2=u^+2as v- final velocity, u- initail velocity as the car stopped after traveling 290m v=0 0=u^2-2*4.3*290 u=sqrt(2494) u=49.94m/s

OpenStudy (anonymous):

Thanks! :)